OpenStudy (anonymous):

Two ropes, AD and BD, are tied to a peg on the ground at point D. The other ends of the ropes are tied to points A and B on a flagpole, as shown below: Two ropes, AD and BD, are tied to a peg on the ground at point D. The other ends of the ropes are tied to points A and B on a flagpole. Angle ADC measures 60 degrees and angle BDC measures 45 degrees. The length of DC is 9. Angle ADC measures 60° and angle BDC measures 45°. What is the distance between the points A and B on the flagpole? 6.59 feet 1.73 feet 4.32 feet 2.56 feet

OpenStudy (anonymous):

@karatechopper

OpenStudy (anonymous):

@Owlcoffee

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@AngusV

OpenStudy (anonymous):

@slade

OpenStudy (anonymous):

2.56 i think @TheSmartOne do u agree

OpenStudy (anonymous):

How do I do it though?

OpenStudy (misssunshinexxoxo):

Which one do you believe it is @CaptainLlama49

OpenStudy (anonymous):

Im not sure how to do it

OpenStudy (misssunshinexxoxo):

This is a isosceles triangle

OpenStudy (anonymous):

Distance formula?

OpenStudy (misssunshinexxoxo):

Locate AC using tan(60) = AC/9 and find BC using tan(45) = BC/9 Then find AB using AB = AC - BC

OpenStudy (anonymous):

2.56?

OpenStudy (misssunshinexxoxo):

Follow this; AB = 9*tan(60)-9*tan(45)

OpenStudy (misssunshinexxoxo):

Plug it in and will give the answer

OpenStudy (misssunshinexxoxo):

Be confident and trust me; I know you could do this! ;)

OpenStudy (anonymous):

-9+9√3

OpenStudy (misssunshinexxoxo):

Make sure if using calculator is in degree mode

OpenStudy (misssunshinexxoxo):

6.59 should be your answer

OpenStudy (anonymous):

It gave me the same thing, something is wrong with my calculator

OpenStudy (misssunshinexxoxo):

Not at all; A is your answer

OpenStudy (anonymous):

Thanks for guiding me through :)